### v1

@@ 4 Phases of compressions and Schur complements
As introduced in Section 1, there is an interlacing relation between the magnitudes of a matrix and those of its submatrices. In this section, we establish its counterpart between the phases of a cramped matrix and those of its compressions and Schur complements.
Let $U$ be an $n\times(n-k)$ isometry, then $\tilde{C}=U^{*}CU$ is said to be a compression of $C$ [3, p. 59]. The phases of $\tilde{C}$ and those of $C$ have the following interlacing property.
Theorem 4.1.
Let $C\in\mathbb{C}^{n\times n}$ be cramped with phases in $(\theta,\theta+\pi)$, $\theta\in[-\pi,\delta(C))$, and $\tilde{C}\!\in\!\mathbb{C}^{(n-k)\times(n-k)}$ be a compression of $C$. Then $\tilde{C}$ is also cramped and
(12) $\phi_{j}(C)\geq\phi_{j}(\tilde{C})\geq\phi_{j+k}(C),\text{ for }j=1,\dots,n-k.$
Proof 4.2.
Since $W(\tilde{C})\subset W(C)$ and $W(C)$ is in an open half plane, it follows that W(C~)W($\tilde${C}) is in the same half plane and thus $\tilde{C}$ is cramped.
Now, from the minimax characterization of phases in (8) and the fact $W(\tilde{C})\subset W(C)$, it follows that
$\phi_{j}(\tilde{C})\leq\phi_{j}(C),$
which is the first inequality in (12). Analogously, the second inequality in (12) follows from the maximin characterization of phases in (8). This completes the proof.
@@
Remark 4.3.
Since any full-rank $X\!\in\!\mathbb{C}^{n\times\!(n-k)}$ can be QR-decomposed as X=QRX\!=\!QR, where $Q\in\mathbb{C}^{n\times(n-k)}$ is an isometry and $R\in\mathbb{C}^{(n-k)\times(n-k)}$ is a nonsingular matrix, it follows from Theorem 4.1 that for any cramped $C\in\mathbb{C}^{n\times n}$ and $j=1,\dots,n-k$,
@@
By exploiting the interlacing property of the phases of compressions of a cramped matrix, we can deduce an interlacing property of the phases of the Schur complements of a cramped matrix. Let $C=\begin{bmatrix}C_{11}&C_{12}\\ C_{21}&C_{22}\end{bmatrix}\in\mathbb{C}^{n\times n}$ be cramped, where $C_{11}\in\mathbb{C}^{k\times k}$. In light of Theorem 4.1, $C_{11}$ is cramped and hence nonsingular.
Theorem 4.4.
Let $C\in\mathbb{C}^{n\times n}$ be cramped with phases in $(\theta,\theta+\pi)$, $\theta\in[-\pi,\delta(C))$. Then $C/\penalty 50_{11}$ is also cramped and
@@
Hence, $C^{-1}$ is cramped. Moreover, if $\phi(C)$ take values in the interval $(\theta,\theta+\pi)$, then $\phi(C^{-1})=[-\phi_{n}(C)\ \dots\ -\phi_{1}(C)]$, taking values in the interval $(-\theta\!-\!\pi,-\theta)$.
We partition $C^{-1}$ into $C^{-1}=\begin{bmatrix}(C^{-1})_{11}&(C^{-1})_{12}\\ (C^{-1})_{21}&(C^{-1})_{22}\end{bmatrix}$. Then $C/\penalty 50_{11}^{-1}=(C^{-1})_{22}$. By Theorem 4.1, $(C^{-1})_{22}$ is cramped and thus, $C/\penalty 50_{11}^{-1}$ is cramped and so is $C/\penalty 50_{11}$. In view of (12), we have
@@
Theorem 4.1 can also be used to construct a simple proof to the following result, which is a special case of [6, Theorem 3.11].
Lemma 4.6.
Let $C\!\in\!\mathbb{C}^{n\times n}$ be cramped with phases in $(\theta,\theta+\pi)$, where $\theta\in[-\pi,\delta(C))$. Then
(13) $\max_{X\in\mathbb{C}^{n\times k}\text{ is full rank}}\sum_{i=1}^{k}\phi_{i}(X^{*}CX)$ $=\sum_{i=1}^{k}\phi_{i}(C),$
(14) $\min_{X\in\mathbb{C}^{n\times k}\text{ is full rank}}\sum_{i=1}^{k}\phi_{i}(X^{*}CX)$ $=\sum_{i=n-k+1}^{n}\phi_{i}(C).$
Proof 4.7.
We prove (13) below. The validity of (14) can be shown similarly.
Application of Theorem 4.1 and Remark 4.3 yields
@@
for any full-rank matrix $X\!\in\!\mathbb{C}^{n\times k}$. Let $C=T^{*}DT$ be a sectoral decomposition. Then defining the full-rank $X=T^{-1}\begin{bmatrix}I_{k}&0\end{bmatrix}^{*}$ gives $\sum_{i=1}^{k}\phi_{i}(X^{*}CX)=\sum_{i=1}^{k}\phi_{i}(C)$.
@@ 5 Compound numerical ranges and numerical ranges of compounds
represents the submatrix of $A$ consisting of rows $i_{1},\dots,i_{k}$ and columns $j_{1},\dots,j_{k}$. It is known that the eigenvalues of the $k$th compound of a square $A$ are products ($k$ at a time) of the eigenvalues of $A$. In particular, $A_{(1)}=A$ and when $A\in\mathbb{C}^{n\times n}$, $A_{(n)}=\det A$. Moreover, if $A$ is Hermitian (resp. unitary), $A_{(k)}$ is also Hermitian (resp. unitary). It also holds that $(AB)_{(k)}=A_{(k)}B_{(k)}$ by the Binet-Cauchy theorem. For more details on compound matrices, we refer to [16, Chapter 19] and [8, Chapter 6].
@@
It follows that $\Lambda_{(k)}(A)\subset W(A_{(k)})$, the numerical range of the compound matrix $A_{(k)}$. Note that even if $A$ is cramped, its compound $A_{(k)}$ is not necessarily cramped, which means we can have the origin lying in the interior of $W(A_{(k)})$ for a cramped $A$. This shows that the numerical range of a compound matrix is not generally suitable for characterizing compound spectra from an angular point of view.
@@
The following lemma generalizes [13, Theorem 1.7.6], which covers the special case of $k=1$.
Lemma 5.1.
Let $A,B\in\mathbb{C}^{n\times n}$ for which $B$ is cramped. Then
@@
The following lemma generalizes [13, Theorem 1.7.8], which covers the special case of $k=1$.
Lemma 5.5.
Let $A,B\in\mathbb{C}^{n\times n}$ for which $B$ is cramped. Then
@@ 6 Phases of matrix product
The phases of an individual complex matrix and their properties have been the focus in Sections 2 to 4. Here, we study the product of two cramped matrices $A$ and $B$.
In view of the majorization result on the magnitudes of the product of two matrices as in (5), it would be desirable to have a phase counterpart in the form of
@@
for cramped matrices $A,B$. We know that this is true when $A,B$ are unitary [17, 21]. Unfortunately, this fails to hold in general, as shown in the following example.
{eg}
Let $A,B$ be $n\times n$ positive definite matrices, then $\phi(A)=\phi(B)=0$. However, $AB$ is in general not a positive definite matrix, and hence $\phi(AB)\neq 0$.
@@
Theorem 6.1.
Let $A,B\in\mathbb{C}^{n\times n}$ be cramped matrices with phases in $(\theta_{1},\theta_{1}+\pi)$ and $(\theta_{2},\theta_{2}+\pi)$, respectively, where $\theta_{1}\in[-\pi,\delta(A))$ and $\theta_{2}\in[-\pi,\delta(B))$. Let $\angle\lambda(AB)$ take values in $(\theta_{1}+\theta_{2},\theta_{1}+\theta_{2}+2\pi)$. Then
(15) $\angle\lambda(AB)\prec\phi(A)+\phi(B).$
Proof 6.2.
@@
This is a special form of the more general Lidskii-Wielandt inequality, which can be found in [4]. While it would be desirable to have a counterpart to this inequality for cramped $A$ and $B$ in the form of
@@ 7 Phases of matrix sum
$\mathcal{C}[\alpha,\beta]=\left\{C\in\mathbb{C}^{n\times n}:\!C\text{ is cramped and }\phi_{1}(C)\!\leq\!\beta,\ \phi_{n}(C)\!\geq\!\alpha\right\}.$
The following theorem can be inferred from the discussions on [23, p. 2]. A proof is presented here for completeness.
Theorem 7.1.
@@
Proof 7.2.
Since $A,B\in\mathcal{C}[\alpha,\beta]$, $\beta-\alpha<\pi$, there exists an open half plane containing both $W(A)$ and $W(B)$. In view of the subadditivity of numerical range, i.e., $W(A+B)\subset W(A)+W(B)$ [13, Property 1.2.7], $W(A+B)$ is contained in the same open half plane. Thus, $A+B$ is cramped. Moreover,
@@
(17) $\leq\max_{x\in\mathbb{C}^{n},\|x\|=1}\max(\angle x^{*}Ax,\angle x^{*}Bx)$
@@
(18) $\geq\min_{x\in\mathbb{C}^{n},\|x\|=1}\min(\angle x^{*}Ax,\angle x^{*}Bx)$
@@
This theorem says that when $\beta-\alpha<\pi$, $\mathcal{C}[\alpha,\beta]$ is a convex cone. A special case is given by $\mathcal{C}[0,0]$, i.e., the cone of positive definite matrices. In general, when $\pi<\beta-\alpha<2\pi$, $\mathcal{C}[\alpha,\beta]$ is a cone, but not a convex cone.
@@ 8 Rank robustness against perturbations
The rank of $I+AB$ plays an important role in the field of systems and control [24]. It is straightforward to see that if $\sigma(A)$ and $\sigma(B)$ are sufficiently small, then $I+AB$ has full rank. If the magnitudes of $A$ are known in advance, a problem well studied [24, 10] in the field of robust control is how large the magnitudes of $B$ need to be before $I+AB$ loses rank by 1, by 2, …, or by $k$. It can be inferred from the Schmidt-Mirsky theorem [19, Theorem 4.18] that for a given matrix $A\in\mathbb{C}^{n\times n}$,
(19) $\inf\left\{\overline{\sigma}(B):\mathrm{rank}(I+AB)\leq n-k\right\}=\sigma_{k}(A)^{-1}.$
@@
$\mathcal{C}_{k}[\alpha]=\left\{C\in\mathbb{C}^{n\times n}:C\text{ is cramped and }{\sum_{i=1}^{k}\phi_{i}(C)\leq\alpha},\sum_{i=n-k+1}^{n}\phi_{i}(C)\geq-\alpha\right\}.$
@@
Theorem 8.1.
Let $A\in\mathbb{C}^{n\times n}\!$ be cramped with phases in $(-\pi,\pi)$ and $k=1,\dots,n$. Then $\mathrm{rank}\,(I+AB)>n-k$ for all $B\in\mathcal{C}_{k}[\alpha],\ \alpha\in[0,k\pi)$, if and only if
@@
Proof 8.2.
First, label the eigenvalues of $AB$ so that $\angle\lambda_{1}(AB)\!\geq\!\angle\lambda_{2}(AB)\!\geq\!\dots\geq\angle\lambda_{n}(AB)$. Since $A$ and $B$ are cramped, $I+AB$ loses rank by $k$ only if
@@
For necessity, suppose to the contraposition that $\sum_{i=1}^{k}\phi_{i}(A)+\alpha\geq k\pi$. We construct below a matrix $B\in\mathcal{C}_{k}[\alpha]$ such that the rank of $I+AB$ is nkn-k. Let $A=T^{*}DT$ be a sectoral decomposition. Define $B=T^{-1}ET^{-*}$, where $E=\mathrm{diag}\{e_{1},e_{2},\dots,e_{n}\}$ satisfies
@@
Clearly, $B$ is cramped. Moreover, $\sum_{i=1}^{k}\phi_{i}(B)=\displaystyle\sum_{i=1}^{k}\angle e_{i}\leq\alpha$, $\sum_{i=n-k+1}^{n}\phi_{i}(B)\geq-\alpha$ and
@@
The necessity of $\sum_{i=n-k+1}^{n}\phi_{i}(A)-\alpha>-k\pi$ can be shown in a similar manner.
The specialization of Theorem 8.1 to the case where $k=1$ is of particular importance in the study of robust control. Specifically, the theorem says that for a cramped matrix $A\in\mathbb{C}^{n\times n}$ with phases in $(-\pi,\pi)$ and $\alpha\in[0,\pi)$, there holds $\mathrm{rank}\,(I+AB)=n$ for all $B\in\mathcal{C}_{1}[\alpha]$ if and only if $\alpha<\min\{\pi-\phi_{1}(A),\pi+\phi_{n}(A)\}$.

### v2

@@ 4 Phases of compressions and Schur complements
There is an interlacing relation between the magnitudes of a matrix $C$ and those of its $(n-k)\times(n-k)$ submatrices [2, Corollary III.1.5]:
Let $C\in\mathbb{C}^{n\times n}$ be partitioned as $C=\begin{bmatrix}C_{11}&C_{12}\\ C_{21}&C_{22}\end{bmatrix}$, where $C_{11}\in\mathbb{C}^{k\times k}$. If $C_{11}$ is nonsingular, then the Schur complement [17] of $C_{11}$ in $C$, denoted by $C/\penalty 50_{11}$, exists and is given by $C/\penalty 50_{11}=C_{22}-C_{21}C_{11}^{-1}C_{12}$. It can be inferred from [17, Corollary 2.3] that $\sigma(C/\penalty 50_{11})$ and $\sigma(C)$ satisfy the following interlacing relation:
In this section, we introduce the interlacing properties between the phases of a sectorial matrix and those of its compressions and Schur complements.
Let $U$ be an $n\times(n-k)$ isometry, then $\tilde{C}=U^{*}CU$ is said to be a compression of $C$. The phases of $\tilde{C}$ and those of $C$ have the following interlacing property [18].
Lemma 4.1
Let $C\in\mathbb{C}^{n\times n}$ be sectorial with phases in $(\theta,\theta+\pi)$, $\theta\in[-\pi,\delta(C))$, and $\tilde{C}\!\in\!\mathbb{C}^{(n-k)\times(n-k)}$ be a compression of $C$. Then $\tilde{C}$ is also sectorial and
$\phi_{j}(C)\geq\phi_{j}(\tilde{C})\geq\phi_{j+k}(C),\text{ for }j=1,\dots,n-k.$ (13)
@@
Remark 4.2
Since any full-rank $X\!\in\!\mathbb{C}^{n\times\!(n-k)}$ can be QR-decomposed as X=QRX\!=\!QR, where $Q\in\mathbb{C}^{n\times(n-k)}$ is an isometry and $R\in\mathbb{C}^{(n-k)\times(n-k)}$ is a nonsingular matrix, it follows from Lemma 4.1 that for any sectorial $C\in\mathbb{C}^{n\times n}$ and $j=1,\dots,n-k$,
@@
By exploiting the interlacing property of the phases of compressions of a sectorial matrix, we can deduce an interlacing property of the phases of the Schur complements of a sectorial matrix. Let $C=\begin{bmatrix}C_{11}&C_{12}\\ C_{21}&C_{22}\end{bmatrix}\in\mathbb{C}^{n\times n}$ be sectorial, where $C_{11}\in\mathbb{C}^{k\times k}$. In light of Lemma 4.1, $C_{11}$ is sectorial and hence nonsingular.
Theorem 4.3
Let $C\in\mathbb{C}^{n\times n}$ be sectorial with phases in $(\theta,\theta+\pi)$, $\theta\in[-\pi,\delta(C))$. Then $C/\penalty 50_{11}$ is also sectorial and
@@
Hence, $C^{-1}$ is sectorial. Moreover, if $\phi(C)$ takes values in the interval $(\theta,\theta+\pi)$, then $\phi(C^{-1})=[-\phi_{n}(C)\ \dots\ -\phi_{1}(C)]$, taking values in the interval $(-\theta\!-\!\pi,-\theta)$.
We partition $C^{-1}$ into $C^{-1}=\begin{bmatrix}(C^{-1})_{11}&(C^{-1})_{12}\\ (C^{-1})_{21}&(C^{-1})_{22}\end{bmatrix}$. Then $C/\penalty 50_{11}^{-1}=(C^{-1})_{22}$. By Lemma 4.1, $(C^{-1})_{22}$ is sectorial and thus, $C/\penalty 50_{11}^{-1}$ is sectorial and so is $C/\penalty 50_{11}$. In view of (13), we have
@@
Lemma 4.1 can also be used to construct a simple proof of the following result, as a special case of [19, Theorem 3.11], which will be used in the subsequent sections.
Lemma 4.4
Let $C\!\in\!\mathbb{C}^{n\times n}$ be sectorial with phases in $(\theta,\theta+\pi)$, where $\theta\in[-\pi,\delta(C))$. Then
$\max_{X\in\mathbb{C}^{n\times k}\text{ is full rank}}\sum_{i=1}^{k}\phi_{i}(X^{*}CX)$ $=\sum_{i=1}^{k}\phi_{i}(C),$ (14)
$\min_{X\in\mathbb{C}^{n\times k}\text{ is full rank}}\sum_{i=1}^{k}\phi_{i}(X^{*}CX)$ $=\sum_{i=n-k+1}^{n}\phi_{i}(C).$ (15)
Proof 3
Application of Lemma 4.1 and Remark 4.2 yields
@@
for any full-rank matrix $X\!\in\!\mathbb{C}^{n\times k}$. Let $C=T^{*}DT$ be a sectorial decomposition. Then defining the full-rank $X=T^{-1}\begin{bmatrix}I_{k}&0\end{bmatrix}^{*}$ gives $\sum_{i=1}^{k}\phi_{i}(X^{*}CX)=\sum_{i=1}^{k}\phi_{i}(C)$, which proves (14). The validity of (15) can be shown similarly.
@@ 5 Compound numerical ranges and numerical ranges of compounds
represents the submatrix of $A$ consisting of rows $i_{1},\dots,i_{k}$ and columns $j_{1},\dots,j_{k}$. It is known that the eigenvalues of the $k$th compound of a square $A$ are products ($k$ at a time) of the eigenvalues of $A$. In particular, $A_{(1)}=A$ and when $A\in\mathbb{C}^{n\times n}$, $A_{(n)}=\det A$. Moreover, if $A$ is Hermitian (resp. unitary), $A_{(k)}$ is also Hermitian (resp. unitary). It also holds that $(AB)_{(k)}=A_{(k)}B_{(k)}$ by the Binet-Cauchy theorem. For more details on compound matrices, we refer to [3, Chapter 19] and [20, Chapter 6].
@@
It follows that $\Lambda_{(k)}(A)\subset W(A_{(k)})$, the numerical range of the compound matrix $A_{(k)}$. Note that even if $A$ is sectorial, its compound $A_{(k)}$ is not necessarily sectorial, which means we can have the origin lying in the interior of $W(A_{(k)})$ for a sectorial $A$. This shows that the numerical range of a compound matrix is not generally suitable for characterizing compound spectra from an angular point of view.
@@
The following lemma generalizes [6, Theorem 1.7.6], which covers the special case of $k=1$.
Lemma 5.1
Let $A,B\in\mathbb{C}^{n\times n}$ for which $B$ is sectorial. Then
@@
The following lemma generalizes [6, Theorem 1.7.8], which covers the special case of $k=1$.
Lemma 5.3
Let $A,B\in\mathbb{C}^{n\times n}$ for which $B$ is sectorial. Then
@@ 6 Phases of matrix product
The phases of an individual complex matrix and their properties have been the focus in Sections 2 to 4. Here, we study the product of two sectorial matrices $A$ and $B$.
In view of the majorization result on the magnitudes of the product of two matrices as in (5), it would be desirable to have a phase counterpart of the form
@@
for sectorial matrices $A,B$. We know that this is true when $A,B$ are unitary [21, 22]. Unfortunately, this fails to hold in general, as shown in the following example.
Example 6.1
Let $A,B$ be $n\times n$ positive definite matrices, then $\phi(A)=\phi(B)=0$. However, $AB$ is in general not positive definite, and hence $\phi(AB)\neq 0$.
@@
Theorem 6.2
Let $A,B\in\mathbb{C}^{n\times n}$ be sectorial matrices with phases in $(\theta_{1},\theta_{1}+\pi)$ and $(\theta_{2},\theta_{2}+\pi)$, respectively, where $\theta_{1}\in[-\pi,\delta(A))$ and $\theta_{2}\in[-\pi,\delta(B))$. Let $\angle\lambda(AB)$ take values in $(\theta_{1}+\theta_{2},\theta_{1}+\theta_{2}+2\pi)$. Then
$\angle\lambda(AB)\prec\phi(A)+\phi(B).$ (16)
Proof 7
@@
This is a special form of the more general Lidskii-Wielandt inequality, which can be found in [23]. While it would be desirable to have a counterpart to this inequality for sectorial $A$ and $B$ in the form of
@@ 7 Phases of matrix sum
$\mathcal{C}[\alpha,\beta]=\left\{C\in\mathbb{C}^{n\times n}:C\text{ is sectorial and }\phi_{1}(C)\!\leq\!\beta,\ \phi_{n}(C)\!\geq\!\alpha\right\}.$
Clearly, $\mathcal{C}[\alpha,\beta]$ is a cone. The following theorem can be inferred from the discussions on [11, p. 2]. A proof is presented here for completeness.
Theorem 7.1
@@
Proof 8
Since $A,B\in\mathcal{C}[\alpha,\beta]$, $\beta-\alpha<\pi$, there exists an open half plane containing both $W(A)$ and $W(B)$. In view of the subadditivity of numerical range [6, Property 1.2.7], i.e., $\subset W(A)+W(B) , W(A+B)W(A+B)$ is contained in the same open half plane. Thus, $A+B$ is sectorial. Moreover,
@@
$\leq\max_{x\in\mathbb{C}^{n},\|x\|=1}\max(\angle x^{*}Ax,\angle x^{*}Bx)$ (18)
@@
$\geq\min_{x\in\mathbb{C}^{n},\|x\|=1}\min(\angle x^{*}Ax,\angle x^{*}Bx)$ (19)
@@
This theorem says that when $\beta-\alpha<\pi$, $\mathcal{C}[\alpha,\beta]$ is a convex cone. A special case is given by $\mathcal{C}[0,0]$, i.e., the cone of positive definite matrices.
@@ 8 Rank robustness against perturbations
The rank of $I+AB$ plays an important role in the field of systems and control [24]. It is straightforward to see that if $\sigma(A)$ and $\sigma(B)$ are sufficiently small, then $I+AB$ has full rank. If the magnitudes of $A$ are known in advance, a problem well studied [24, 25] in the field of robust control is how large the magnitudes of $B$ need to be before $I+AB$ loses rank by 1, by 2, …, or by $k$. It can be inferred from the Schmidt-Mirsky theorem [26, Theorem 4.18] that for a given matrix $A\in\mathbb{C}^{n\times n}$,
$\inf\left\{\overline{\sigma}(B):\mathrm{rank}(I+AB)\leq n-k\right\}=\sigma_{k}(A)^{-1}.$ (20)
@@
$\mathcal{C}_{k}[\alpha]=\left\{C\in\mathbb{C}^{n\times n}:C\text{ is sectorial and }{\sum_{i=1}^{k}\phi_{i}(C)\leq\alpha},\sum_{i=n-k+1}^{n}\phi_{i}(C)\geq-\alpha\right\}.$
@@
Theorem 8.1
Let $A\in\mathbb{C}^{n\times n}\!$ be sectorial with phases in $(-\pi,\pi)$ and $k=1,\dots,n$. Then $\mathrm{rank}\,(I+AB)>n-k$ for all $B\in\mathcal{C}_{k}[\alpha],\ \alpha\in[0,k\pi)$, if and only if
@@
Proof 9
First, label the eigenvalues of $AB$ so that $\angle\lambda_{1}(AB)\!\geq\!\angle\lambda_{2}(AB)\!\geq\!\dots\geq\angle\lambda_{n}(AB)$. Since $A$ and $B$ are sectorial, $I+AB$ loses rank by $k$ only if
@@
For necessity, suppose to the contraposition that $\sum_{i=1}^{k}\phi_{i}(A)+\alpha\geq k\pi$. We construct below a matrix $B\in\mathcal{C}_{k}[\alpha]$ such that the rank of $I+AB$ is nkn-k. Let $A=T^{*}DT$ be a sectorial decomposition. Define $B=T^{-1}ET^{-*}$, where $E=\mathrm{diag}\{e_{1},e_{2},\dots,e_{n}\}$ satisfies
@@
Clearly, $B$ is sectorial. Moreover, $\sum_{i=1}^{k}\phi_{i}(B)=\displaystyle\sum_{i=1}^{k}\angle e_{i}\leq\alpha$, $\sum_{i=n-k+1}^{n}\phi_{i}(B)\geq-\alpha$ and
@@
The necessity of $\sum_{i=n-k+1}^{n}\phi_{i}(A)-\alpha>-k\pi$ can be shown similarly.
The specialization of Theorem 8.1 to the case where $k=1$ is of particular importance in the study of robust control. Specifically, the theorem says that for a sectorial matrix $A\in\mathbb{C}^{n\times n}$ with phases in $(-\pi,\pi)$ and $\alpha\in[0,\pi)$, there holds $\mathrm{rank}\,(I+AB)=n$ for all $B\in\mathcal{C}_{1}[\alpha]$ if and only if $\alpha<\min\{\pi-\phi_{1}(A),\pi+\phi_{n}(A)\}$.